OpenStudy (anonymous):

Find the gradient of the curve x^2 + 2xy-2y^2 +x=2 at the point (-4,1)

OpenStudy (anonymous):

First, you find out partial x and partial y.. do you have that yet?

OpenStudy (fibonaccichick666):

I'd suggest starting with what is a gradient then go into partial derivatives

OpenStudy (anonymous):

Does this help?\[ x^2~+~2xy~-~2y^2~+x~=2 \\ 2x+2\frac{\textrm{d}y}{\textrm{d}x}-4\frac{\textrm{d}y}{\textrm{d}x}+1~=~0 \]I can't remember if gradient is the same thing as slope.

OpenStudy (amistre64):

gradient means slope in certain context yes; its also a term used to find the normal to a surface

OpenStudy (amistre64):

i think you missed your product rule tho

OpenStudy (amistre64):

\[x^2~+~2xy~-~2y^2~+x~=2 \] \[2xx'~+~2x'y+~2xy'~-~4yy'~+x'~=0 \] if "wrt.x" then x' = dx/dx = 1 \[2x~+~2y+~2xy'~-~4yy'~+1~=0 \]

OpenStudy (amistre64):

plug in (-4,1) and solve for y'